# Choose Your Diffusion: Efficient and flexible ways to accelerate the diffusion model in fast high energy physics simulation

### v1

@@ 5.1 CaloChallenge
Figure 8: Separation power for $E_{\mathrm{ratio}}$ with different samplers and schedulers as a function of sampling steps, the shaded area is possible values of separation power for 20 independent samplings.
Figure 9: The distribution for total shower energy divided by the incident particle energy by left: 400 steps DDPM, middle: 30 steps Restart samplers, left: high step EDM samplers
Samplers(steps) AUC Separation power FPD
DDPM(79) 55.3 0.0810 0.074
DDPM(200) 53.2 0.0344 0.046
DDPM(400) 52.6 0.0155 0.043
EDM(36) 54.1 0.0256 0.035
EDM(79) 52.0 0.0076 0.027
EDM(200) 51.5 0.0055 0.023
EDM(400) 51.1 0.0039 0.021
EDM_DPM++(79) 52.3 0.0103 0.026
EDM_Lu(79) 52.2 0.0086 0.026
Restart(18) 55.2 0.0169 0.059
Restart(36) 52.0 0.0073 0.025
Restart(79) 51.8 0.0057 0.022
LMS(36) 53.8 0.0305 0.095
DPM++(20) 59.8 0.0534 0.146
UniPC(20) 60.3 0.1304 0.152
@@
Table 1: Summary table for high-level feature classifier mean AUC, $E_{\mathrm{ratio}}$ separation power, and Frechét Particle Distance [52] of of different samplers. (400-step DDPM as baseline)
@@
Now we take a look at one of the most difficult and important quantities for the low level voxel diffusion – $E_{\mathrm{ratio}}$. In Fig.8, we choose 6 cases in our study. First, much faster convergences have been seen from all new introduced samplers by 5-20 times. The result matches with the Table.3 mentioned in Song’s [53]. For EDM Heun SDE sampler, the separation power quickly drops around 60 steps and achieves better performance than 400 steps DDPM. If we zoom-in the figure, at around 100-200 steps, the separation power becomes 3-5 times weaker than the baseline DDPM samplers. In Fig. 9, we have seen really aligned shape between generated and reference classes for Restart and EDM samplers. As the steps goes even higher, the discrepancy between two histogram become smaller while the shape remains aligned. The histogram geneated from Restart samplers does not fit perfectly on lower ratio region, but still better than DDPM in other regions. This is crucial for us to perform accurate energy calibration from low-level fast calorimeter simulation later.
@@
The preliminary Frechét Particle Distance (FPD) [52] values for some samplers are also provided in Table. 1. The FPD metric is relatively biased and unstable among sub samples we choose. So we follow the way from [20] to subtract the non-zero FPD value between two GEANT4 sample. Surprisingly, we find the EDM sampler converges even faster in this metric, surpassing the DDPM sampler around 36 steps. The 18-step Restart sampler has slightly worse performance, this is mainly because we choose the predefined configuration for smallest $E_{\mathrm{ratio}}$ separation power, a throughout fine-tune can make FPD value 20-30 percent smaller with almost same generation time. However, it still obtains the best FPD results with larger steps.
@@
We choose Karras and Lu schedulers to illustrate the impacts of different noise schedulers on the same samplers. Changing the Karras scheduler to Lu will decrease the separation power in smaller steps ($\leq 100$) but increase it in larger steps. The choice of schedulers provides a training-free method to balance and control the quality and speed of the backward diffusion process. The EDM/DPM-Solver combined sampler with Karras scheduler performs pretty similarly as the EDM sampler with Lu scheduler. The best performances of Restart samplers again are achieved in 30-36 steps. It is crafted for efficiently conducting fast ODE sampling with outstanding performance and can be fine-tuned with post-training configuration. The DPM-solver exhibits a rapid drop in separation power within 10-20 steps but ceases to improve beyond those steps, achieving better performance than the 100-step DDPM around the 15th step. The plot further confirms that ODE sampler does better in low sampling steps as the discretization error of them are smaller than SDE ones in this region, the contraction of accumulated errors from SDE sampling process become much stronger than ODE so that a better convergence observed in higher sampling steps than ODE ones. Some ODE samplers, such as LMS and DPM-solver, perform well on some metrics but not all. There is not a clearly superior option that can achieve the best quality while having the lowest generation time.
@@
Figure 12: Comparison plots for relative jet mass with 200 steps DDIM Euler Sampler, 18 steps multi-level Restart Sampler, 50 steps EDM samplers . left: gluon initiated jet, right: top quark jet.
Samplers(steps) AUC (Gluon Jet) FPD (Gluon Jet) AUC (Top Jet) FPD (Top Jet)
Euler(200) 52.6 0.15 53.5 0.19
EDM(50) 51.7 0.05 51.9 0.07
Restart(18) 51.6 0.04 51.5 0.05
Restart_multi(18) 51.2 0.02 51.2 0.02
Table 2: Summary table for low-level feature classifier mean AUC and Frechét Particle Distance calculated from jet-level classifier of different samplers on gluon-initiated jets top quark jets
@@ 5.2 JetNet
We employ the same network training techniques and representative samplers for the JetNet dataset to demonstrate that these methods are entirely portable across various datasets. As depicted in Fig. 1011, the low level feature generations with a 50-step EDM sampler and 18-step Restart sampler show better performance than a 200-step DDIM [33] Euler sampler on gluon jets. For more complicated topology, EDM and Restart have advantages in simulating the tails of the distribution over large $\eta$ region. EDM sampler has small deficiency for underestimating the particles from top quark jets after $\Delta\phi>0.2$. This would have impact on reconstructing the jet level variables as shown in Fig. 12. When simulating the complex distribution in a highly boosted environment, the EDM and DDIM Euler samplers fail to achieve satisfactory alignment around the two asymmetric peaks in the relative jet mass distribution. The low step Restart sampler gives obviously better generated distribution than others. We evaluate the samplers via the three low level features with the same DNN classifier, and the values for the Fréchet Particle distance (FPD) calculated from the Particle [54] multi-classifier with N = 6. Performances are summarized on Table. 2.
The boosts brought by those changes isn’t as pronounced as they are in CaloChallenge. It may be because methods are more applicable to UNet and pixelated data than point clouds network.

### v2

@@ 5.1 CaloChallenge
Figure 8: Separation power for $E_{\mathrm{ratio}}$ with different samplers and schedulers as a function of evaluation steps, the shaded area is possible values of separation power for 20 independent samplings.
We showed the possible AUC values for high level feature classifiers with different samplers and schedulers by evaluating different samplers 20 times in Fig. 7. DPM-solver sampler seems to perform poorly on predicting the low and high-level features simultaneously. The Uni-PC sampler has very similar performance with DPM-solver sampler and only gives slightly better performance over very low sampling (10-20) steps. The lowest AUC is around 57 at 50th step. The EDM and Restart samplers surpass the previous 400 steps DDPM samplers at 50, 25 steps (100, 60 function evaluations) correspondingly, combined with other plots we showed, achieving the same and better performance with a 3-8 times speedup. The LMS ODE sampler has a better performance than EDM below 45 steps, but reaches a worse convergence than other samplers. The best AUC value for Restart samplers achieve at 30-36 and 79 steps, this might because we use the predefined configuration specifically for those steps. All proposed samplers have comparable and better performance with smaller steps. Most of the samplers we use actually have more parameters to tune than the baseline DDPM. For the common ODE samplers, a closer matching noise level during sampling would often yield better results. A possible future development is to fine-tune the scheduler configuration for ODE samplers in higher steps.
Samplers(steps) AUC $\downarrow$ Separation power $(10^{-2})$ $\downarrow$ FPD $\downarrow$
DDPM(79) $\pm$0.05 $\pm$1.05 $\pm$0.01
DDPM(200) $\pm$0.03 $\pm$0.40 $\pm$0.007
DDPM(400) $\pm$0.02 $\pm$0.30 $\pm$0.007
EDM(36) $\pm$0.06 $\pm$0.72 $\pm$0.007
EDM(79) $\pm$0.04 $\pm$0.20 $\pm$0.004
EDM(200) $\pm$0.03 $\pm$0.10 $\pm$0.004
EDM(400) $\pm$0.01* $\pm$0.05* $\pm$0.003*
EDM_DPM++(79) $\pm$0.04 $\pm$0.20 $\pm$0.004
EDM_Lu(79) $\pm$0.04 $\pm$0.20 $\pm$0.004
Restart(18) $\pm$0.07 $\pm$0.66 $\pm$0.009
Restart(36) $\pm$0.06 $\pm$0.26 $\pm$0.003
Restart(79) $\pm$0.05 $\pm$0.19 $\pm$0.004
LMS(36) $\pm$0.04 $\pm$1.35 $\pm$0.01
DPM++(25) $\pm$0.07 $\pm$0.51 $\pm$0.01
UniPC(25) $\pm$0.07 $\pm$0.45 $\pm$0.01
@@
Table 1: Summary table for high-level feature classifier mean AUC, $E_{\mathrm{ratio}}$ separation power, and Frechét Particle Distance [58] of of different samplers. Underline is the 400-step DDPM sampler (CaloDiffusion) as baseline, the *, †, ‡, are the first, second, third best results. Bold is the highlighted result with either great performance or good quality-sampling time balance.
@@
Now we take a look at one of the most difficult and important quantities for the low level voxel diffusion – $E_{\mathrm{ratio}}$. In Fig.8, we choose 6 cases in our study. First, much faster convergences have been seen from new introduced samplers by 5-20 times. The obtained result matches with the Table.3 mentioned in Song’s [59]. For EDM sampler, the separation power, defined as the triangular discriminator [8] $\frac{\sum_{i=1}^{n}\left(\text{count}_{\text{data},i}-\text{count}_{\text{gen},i}\right)^{2}}{\sum_{i=1}^{n}\left(\text{count}_{\text{data},i}+\text{count}_{\text{gen},i}\right)}$ which calculated from each bins of the histogram, quickly drops around 60 steps and achieves better performance than the baseline 400 steps DDPM sampler. The 20-30 steps DPM-Solver and UniPC samplers already have lower separation powers than 200 steps DDPM sampler. The 18 steps Restart sampler already could match the similar performance as the baseline. If we zoom-in the figure, at around 100-200 steps, EDM and Restart samplers have 3-5 times weaker separation power than the baseline DDPM samplers.
@@
The preliminary Frechét Particle Distance (FPD) [58] values for some samplers are also provided in Table. 1. The FPD metric is relatively biased and unstable among subsamples we choose. So we follow the way from [23] to subtract the non-zero FPD value between two GEANT4 sample. Surprisingly, we find the EDM sampler converges even faster in this metric, surpassing the DDPM sampler around 36 steps (72 NFE). The 18-steps ($\sim$ 40 NFE) Restart sampler has slightly worse performance, this is mainly because we choose the predefined configuration for smallest $E_{\mathrm{ratio}}$ separation power, a throughout fine-tune can make FPD value 20-30 percent smaller with almost same generation time. However, it still obtains the best FPD results with larger steps.
@@
The separation power is decreased in smaller steps ($\leq 100$) but increased in larger steps by changing the EDM scheduler from Karras to Lu. The choice of schedulers provides a training-free method to balance the quality and speed of the backward diffusion process. The EDM/DPM-Solver combined sampler with Karras scheduler performs pretty similarly as the EDM sampler with Lu scheduler. The best performances of Restart samplers again are achieved in 30-36 steps. This sampler is crafted for efficiently conducting fast generation with outstanding performance and can be fine-tuned with post-training configuration. The DPM-solver exhibits a rapid drop in separation power within 10-20 steps but ceases to improve beyond those steps, achieving better performance than the 100-step DDPM around the 15th step. The plot further confirms that ODE sampler does better in low sampling steps as the discretization error of them are smaller than SDE ones in this region, the contraction of accumulated errors from SDE sampling process become much stronger than ODE so that a better convergence observed in higher sampling steps than ODE ones [48]. Some ODE samplers, such as LMS and DPM-solver, perform well on some metrics but not all. There is not a clearly superior option that can achieve the best quality while having the lowest generation time.
@@
Figure 11: Comparison plots for substructure of particle constituents inside gluon/top quark jets with different samplers and schedulers.
Samplers(NFE) AUC $\downarrow$ FPD $\downarrow$ $W_{m}(10^{-4})$ $W_{p_{T}}(10^{-4})$ $W_{\tau_{32}}(10^{-3})$ $\downarrow$
Top
DDPM(200) $\pm$0.04 $\pm$0.03 13.54 12.04 16.05
Euler(200) $\pm$0.04 $\pm$0.03 20.47 10.08 11.54
EDM(100) $\pm$0.03 $\pm$0.01 7.65 7.07 8.32
EDM_DPM++(100) $\pm$0.02 $\pm$0.005 8.05 7.17 8.34
Restart(70) $\pm$0.03* $\pm$0.005* 4.30* 5.65* 5.90*
DPM++(25) $\pm$0.05 $\pm$0.04 25.64 10.45 25.76
UniPC(25) $\pm$0.07 $\pm$0.03 27.89 10.49 26.04
Gluon
DDPM(200) $\pm$0.04 $\pm$0.03 5.69 6.01 14.30
Euler(200) $\pm$0.04 $\pm$0.04 6.40 6.63 4.79
EDM(100) $\pm$0.03 $\pm$0.01 5.03 4.65 4.92
EDM_DPM++(100) $\pm$0.03 $\pm$0.01 5.40 4.59 5.03
Restart(70) $\pm$0.02* $\pm$0.007* 3.03* 3.20* 3.15*
DPM++(25) $\pm$0.05 $\pm$0.02 6.05 9.97 10.09
UniPC(25) $\pm$0.05 $\pm$0.02 6.15 9.98 5.41
Table 2: Summary table for low-level feature classifier mean AUC, Frechét Particle Distance calculated from jet-level classifier, and high-level features unbinned (Wasserstein) distance of different samplers on gluon-initiated jets top and gluon quark jets. Underline is the 400-step DDPM sampler (PC-JeDi) as baseline, the *, †, ‡, are the first, second, third best results. Bold is the highlighted result with either great performance or good quality-sampling time balance.
Figure 12: Comparison plots for FPD values for top jets as a function of evaluation steps with different samplers and schedulers. The shaded area is possible values of separation power for 10 independent samplings.
Figure 13: Comparison plots for FPD values for gluon jet as a function of evaluation steps with different samplers and schedulers. the shaded area is possible values of separation power for 10 independent samplings.
@@ 5.2 JetNet
We employ network training techniques with the same continuous noise level and representative samplers for the JetNet dataset to demonstrate that these methods are entirely portable across various datasets. As depicted in Fig. 9 and 10, the low level feature generations with a 50-steps EDM sampler and 18-steps Restart sampler show better performance than a 200-steps DDIM [36] Euler sampler on gluon jets. For more complicated topology, EDM and Restart have advantages in simulating the tails of the distribution over large $\eta$ region. EDM sampler has small deficiency for underestimating the particles from top quark jets after $\Delta\phi>0.2$. This would have impact on reconstructing the jet level variables in Fig. 11.
The same training setup in PC-JeDi [25] is used with DDPM samplers as baseline. We found the Euler samplers could give a better performance with our new proposed training. Both EDM and Restart samplers capture most of the high level gluon jet features, with only few percent difference in relative transverse momentum and mass of jets. The improvement is clearer in more challenging substructure variables such as the n-subjetiness between subleading and third leading jets - $\tau_{32}$ when comparing with default DDPM and Restart samplers. Although some ODE samplers outperform than DDPM on $\tau_{32}$ of gluon jets, they can not achieve better overall performance, except LMS samplers. The largest discrepancy for this quantity appears around 0.7-0.9 region where the baseline DDPM sampler has around 30-40% difference. The low-steps Restart sampler gives obviously better generated distribution than others with less than 10% difference over most regions.
While simulating the complex distribution in a highly boosted environment, most of the samplers fail to achieve satisfactory alignment around the two asymmetric peaks in the relative mass distribution for top jets. We then anatomise the performance of proposed samplers via the three low level features with the same DNN classifier used in previous section, the values for the Fréchet Particle distance (FPD) calculated from the ParticleNet [60] multi-classifier with N = 6, and the unbinned metrics (Wasserstein distance) for reconstructed individual high level features. Performances are summarized on Table. 2. The Restart sampler provides the consistently best results for both top and gluon jets in all three metrics. Different scheduler variants on SDE samplers should not have too much influence on results. The 25-steps DPM-Solver sampler shows competitive result as the 200-steps traditional Euler and DDPM samplers on gluon jets.
In Fig. 1213, Restart sampler yields the best performance even in small number of function evaluations. Even though the boosts brought by those changes isn’t as pronounced as they are in CaloChallenge, high step EDM and Restart samplers still have much lower FPD values than other samplers.